OpenStudy (anonymous):

tuition was $2800. a loan was for 6/7of tuition . how much was the loan

OpenStudy (anonymous):

2800/7 = 400 = 1/7 400 * 4 = 1600 = 4/7 The loan was $1600

OpenStudy (anonymous):

(l2800*6)/7

OpenStudy (texaschic101):

2800 x (6/7)